# Programming Quantum Measurements of Time inside a Complex Medium

### v1

@@ I Introduction
Here we show how the coupling of spatial and temporal information inside an MMF can be harnessed to program generalized measurements for large photonic time-bin qudits. In particular, we use the fiber MSTM to construct a special set of spatial modes that arrive through the MMF unscattered in time. By exciting arbitrary coherent linear superpositions of these modes via wavefront shaping techniques, we are able to engineer large, multi-mode, unbalanced interferometers inside the MMF without the need for any phase stabilization. In this manner, the MMF functions as a generalized projective measurement in the time-bin basis, allowing us to perform time-bin measurements in mutually unbiased bases (MUBs) in dimensions up to $d=11$ inside a 40m-long commercial step-index MMF without any phase stabilizer. We certify the quality of our measurements by performing quantum tomography via a prepare-and-measure scheme that uses a tuneable Franson interferometer to prepare arbitrary time-bin qubits in all two-dimensional subspaces of our high-dimensional measurement Hilbert space. Our fiber-based method for realizing generalized, high-dimensional measurements of photonic time-bin can readily be scaled to larger time-bin dimensions while not requiring any active phase stabilization, making it particularly exciting for many quantum technology applications. Below, we present the central concept behind our technique, discuss its experimental implementation, and outline its advantages and current limitations.
@@ Acknowledgements.
We would like to acknowledge David Phillips for helpful discussions and extend our thanks to Joel Carpenter for his library “digHolo” for performing high-speed off-axis holography. We acknowledge financial support from the UK Engineering and Physical Sciences Research Council (EPSRC) (EP/Z533208/1), European Research Council (ERC) Starting Grant PIQUaNT (950402), and the Royal Academy of Engineering Chair in Emerging Technologies programme (CiET-2223-112).

### v2

@@ I Introduction
Here we show how the coupling of spatial and temporal information inside an MMF can be harnessed to program generalized measurements for large photonic time-bin qudits. In particular, we use the fiber MSTM to construct a special set of spatial modes that arrive through the MMF unscattered in time. By exciting arbitrary coherent linear superpositions of these modes via wavefront shaping techniques, we are able to engineer large, multi-mode, unbalanced interferometers inside the MMF. In this manner, the MMF functions as a generalized projective measurement in the time-bin basis, allowing us to perform time-bin measurements in mutually unbiased bases (MUBs) in dimensions up to $d=11$ inside a 40m-long commercial step-index MMF. We certify the quality of our measurements by performing quantum tomography via a prepare-and-measure scheme that uses a tuneable Franson interferometer to prepare arbitrary time-bin qubits in all two-dimensional subspaces of our high-dimensional measurement Hilbert space. Our fiber-based method for realizing generalized, high-dimensional measurements of photonic time-bin can readily be scaled to larger time-bin dimensions while not requiring any active phase stabilization, making it particularly exciting for many quantum technology applications. Below, we present the central concept behind our technique, discuss its experimental implementation, and outline its advantages and current limitations.
@@ Acknowledgements.
We would like to acknowledge David Phillips for helpful discussions and extend our thanks to Joel Carpenter for his library “digHolo” for performing high-speed off-axis holography. We acknowledge financial support from the UK Engineering and Physical Sciences Research Council (EPSRC) (EP/Z533208/1) (EP/Z533166/1), European Research Council (ERC) Starting Grant PIQUaNT (950402), and the Royal Academy of Engineering Chair in Emerging Technologies programme (CiET-2223-112).